OpenStudy (perl):

Johnny has been planting blueberries in a garden in his backyard. Johnny started with blueberries in one square foot of garden modeled by the function s(x) = 10. He is allowed to convert more of the garden, and each additional square foot he maintains allows his blueberry plants to produce more by a rate of a(x) = 1.4^(x-2). Explain to Johnny how to create an equation to predict the number of blueberries he can expect based on the number of square feet he maintains. Describe how to determine the number of blueberries he will grow with 5 square feet

OpenStudy (perl):

I am stumped on the part that says 'more by a rate of '

OpenStudy (perl):

i was asked to help another student with this, and i find the wording a bit odd

OpenStudy (zzr0ck3r):

and this is not calc?

OpenStudy (perl):

algebra *2

OpenStudy (perl):

I tried this y = 10( 1 + 1.4^(x-2)) , but i dont think thats right, for x>= 1

OpenStudy (zzr0ck3r):

nm yeah I dont know this is confusing lol

OpenStudy (perl):

:D tell me about it

OpenStudy (zzr0ck3r):

it seems like we need to multiply the rate times some amount of time, so that our units are berries

OpenStudy (zzr0ck3r):

what does f(x) tell us? the amount of berries?

OpenStudy (perl):

correct

OpenStudy (zzr0ck3r):

lol

OpenStudy (perl):

you found it :)

OpenStudy (perl):

i suspect the question might have a typo , but in the worst case scenario

OpenStudy (zzr0ck3r):

it might be s(x) * a(x)

OpenStudy (zzr0ck3r):

but that does not work with the units....

OpenStudy (zzr0ck3r):

\(b * \frac{b}{t} \ne b\)

OpenStudy (zzr0ck3r):

I dont know, but if you figure it out...let me know.

OpenStudy (perl):

also, here is another question

OpenStudy (perl):

A field test for a new exam was given to randomly selected seniors. The exams were graded, and the sample mean and sample standard deviation were calculated. Based on the results, the exam creator claims that on the same exam, nine times out of ten, seniors will have an average score within 4% of 70%. Is the confidence interval at 90%, 95%, or 99%? What is the margin of error? Calculate the confidence interval and explain what it means in terms of the situation.

OpenStudy (perl):

hey, how would you prove rigorously that f(x,y) = x^2 - y^2 is onto, on R? just wondering